OpenStudy (anonymous):

an open tube mercury manometer has a height difference between the columns of -10cm. What are the a) gauge pressure and b) the absolute pressure?

OpenStudy (anonymous):

Gamma*h= change in pressure, So what's the specific weight of mercury and times that by .1meters

OpenStudy (anonymous):

whats the formula for gamma?

OpenStudy (anonymous):

Gamma is specific weight, look it up in a table. It will be a number with the units of N/m^3

OpenStudy (anonymous):

ahhh,,,after getting the pressure? and multipying it by .1 meters? i will now get its gauge and absolute? right?

OpenStudy (anonymous):

the specific weight for mercury is 133416 N/m3

OpenStudy (anonymous):

Mmm... I probably didn't make it clear before: Gauge pressure = specific weight of mercury * change in height in meters Absolute pressure = gauge pressure + atmospheric pressure in pascals

OpenStudy (anonymous):

multipied by .1m = 13341.6

OpenStudy (anonymous):

it will goe like this,, gauge pressure = 133416 N/m3 * -.1 m = -13341.6 Pa absolute pressure = -13341.6 + 101325 Pa = 87983.4 Pa

OpenStudy (anonymous):

*goes

OpenStudy (anonymous):

can it be negative pa in gauge?

OpenStudy (anonymous):

Looks like u did it right ^_^

OpenStudy (anonymous):

can it be negative pa in gauge?